# On the bouncing completion of eternal inflation

### v1

@@ I Introduction
This was perceived by Borde, Guth and Vilenkin, who contrived it into the BGV theorem N3 ; Borde_1994 , providing a route to verify incompleteness: if the average expansion of the spatial sections is positive, the geometry is geodesically incomplete.
@@
Recently, the theorem gained renewed attention as cyclic cosmologies IS19 ; IS21 ; P10 were shown to be incomplete KS22 ; N4 . Further, counter-examples were found due to the existence of loopholes in the construction of the average expansion LED23 , afterwards amended in a re-proposal GHZ24 . Still, past eternal inflation must be modified and, together with any eternal Universe proposal, can only be complete at the cost of breaking the null energy condition EL24a ; EL24b ; carballorubio2024 . However, a fundamental assumption in all approaches is that the space-time geometry is inextensible N2 ; Wald1984rg ; Penrose65 ; GEROCH1968526 ; Hawking66-I ; Hawking66-II ; Hawking67-III ; Hawking-Penrose ; Ellis1977 ; Senovilla_2015 . Thus, the question of whether an averaged expanding space-time can be past eternal depends on whether it admits an extension or not, which enters as an hypothesis in the BGV theorem. Along this route, extendibility conditions for singular space-times were pursued in Clarke-local-extensions ; Clarke-local-extensionsII ; Clarke-singularities-hyperbolicST , and for flat FLRW in Klein_2017 ; LP17 ; YQ18 ; HISC22 ; Nomura_2021 , through co-moving coordinates, curvature tensors in parallelly propagated tetrads and/or null-like coordinates for the case of quasi-de Sitter N5 . The aim of this paper is to propose an alternative extension of such space-times through the construction of a physically motivated chart given by the non co-moving observer reference frame. As an intuitive argument note that, together with the infinite time dilation of the co-moving observer clock, there must be an infinite contraction of their spatial rulers, inducing a zero spatial volume in this asymptotic limit. Hence, the spatial volume seen by the non co-moving observers in this asymptotic limit may be finite, allowing its extension. Indeed, this approach allows us to exhaust all possible cases in flat FLRW space-times, showing that only past asymptotic de Sitter (dS) can be extended (excluding the cyclic models presented in Refs. IS19 ; IS21 ). There, the non co-moving observer sees his co-moving volume reaching a velocity dependent finite minimum, before growing again: the extension must contain a bounce. This paper is organized as follows: In Section II we revisit the BGV theorem, restating its assumptions, some loopholes, and the revised proposal. In section III we present a local analytic extension to non co-moving observers in the spatially flat dS space-time. In section IV we derive necessary conditions for a $\mathcal{C}^{2}$ metric extension of an arbitrary spatially flat FLRW space. We conclude in section V.
@@ II The Borde-Guth-Vilenkin Theorem
with geodesic equations:
@@
$H_{BGV}(s)=\frac{-u^{\mu}_{0}}{\sigma^{2}+\kappa}\frac{Du_{\mu}}{\textrm{d}s}=\frac{-1}{\sigma^{2}+\kappa}\frac{\textrm{d}\sigma}{\textrm{d}s}=-\frac{\textrm{d}}{\textrm{d}s}F(\sigma),$ (3)
being $\sigma=u_{0}^{\nu}u_{\nu}$ and $\kappa=u_{0}^{\nu}u_{0\nu}$. For a time-like $u^{\mu}_{0}$, $\sigma$ represents the relative Lorentz factor and $\kappa=-1$, while for a light-like $u^{\mu}_{0}$, $\sigma=\textrm{d}t/\textrm{d}s$, the change of time w.r.t. the affine parameter, and $\kappa=0$. Moreover,
$F(\sigma)=\begin{cases}\sigma^{-1}&\text{when $\kappa=0$,}\\ \frac{1}{2}\mathrm{ln}\frac{\sigma+1}{\sigma-1}&\text{when $\kappa=-1$.}\\ \end{cases}$ (4)
Notice that the generalized Hubble parameter reduces to the usual one $H=(1/3)u^{\mu};_{\mu}$ when smeared over all directions of $u^{\mu}_{0}$ N3 . Further, defining the averaged expansion rate $\mathcal{H}_{avg}$ as
@@
$\int^{s_{f}}_{s_{i}}H_{BGV}\,\textrm{d}s=F(\sigma_{f})-F(\sigma_{i})\leq F(\sigma_{f}).$ (6)
@@
$0<\mathcal{H}_{avg}\leq\frac{F(\sigma_{f})}{\Delta s},$ (7)
@@
Nonetheless, the present work intends to scrutinize a different aspect of the BGV theorem, namely, the assumption that the space-time is inextensible. Important works have been done in this direction Klein_2017 ; LP17 ; YQ18 ; HISC22 ; Nomura_2021 . However, here we seek a different approach by constructing a chart adapted to the incomplete non co-moving observer, allowing us to grasp some physical intuition on possible extensions and the elusive nature of some singular geometries.
@@
The flat coordinates in dS space-time is a pivotal example. Given by eq. (1) with an exponential scale factor $a(t)=a_{dS}(t)=\textrm{e}^{\alpha t}$ it satisfies all the BGV theorem hypothesis and is therefore incomplete, albeit the absence of any singularity, indicating that the incompleteness is just a limitation of the flat patch.
@@
$r$ $=l-\frac{1}{\alpha v_{0}}\left[\mbox{cotgh}(\Theta-\Theta_{0})+\mbox{cotgh}\Theta_{0}\right],$ (11)
@@
Equations (10,11) give the coordinate transformations from the co-moving coordinates to the non co-moving ones. Direct derivation w.r.t. the parameter $l$ leads to the normalized vector field $v^{\mu}=(v^{t},v^{r},0,0)$ tangent to the space-like hyper-surfaces in the new coordinate system:
@@
where $a\equiv a_{dS}(\tau,l)$ and $r\equiv r(\tau,l)$ as in Eqs. (10) and (11).
Although $a_{dS}$ is expressed in terms of a hyperbolic sine, it depends on $l$ also, and the metric above is spatially inhomogeneous and anisotropic. Hence, it cannot be understood as the incomplete dS with negative spatial curvature with a zero volume element when $a=0$, as it was done in Ref. N4 , which is clearly not the case here. Note that $l$ is not a radial coordinate.
@@
which reaches a minimum $\sqrt{-g}=|v_{0}|\mbox{sin}\theta/\alpha^{2}$ when $\Theta=\Theta_{0}$, growing again for $\Theta<\Theta_{0}$. This is expected as the dS manifold can be covered by two flat patches, and if we assume that time for the non co-moving observer goes only in one direction, one sheet is contracting while the other expands. This is in contrast with Ref. N6 , in which the complete dS space-time is covered by two expanding sheets, and time runs in different directions and no flow of energy between the sheets must be imposed as a boundary condition. The argument is that the energy density of non co-moving observers, the $T^{0}_{0}$ component of its energy-momentum tensor, diverges in the boundary. However, this is only a coordinate problem. For the co-moving observer the non co-moving one approaches light velocity near this boundary, appearing to have infinity energy, but nothing happens from the point of view of non co-moving observers themselves, analogous to what happens near a black hole’s horizon, where observers can naturally pass through it from their own point of view.
@@
$u^{\mu}_{;\mu}=\alpha\mbox{tanh}(\Theta-\Theta_{0})+\\ 2\alpha\left[\frac{(\alpha v_{0}l+\mbox{cotgh}\Theta_{0})\mbox{cotgh}(\Theta-\Theta_{0})-1}{\alpha v_{0}l-\mbox{cotgh}(\Theta-\Theta_{0})+\mbox{cotgh}\Theta_{0}}\right].$ (15)
When $\Theta=\Theta_{0}$ the expansion becomes $\theta=2\alpha(\alpha v_{0}l+\mbox{cotgh}\Theta_{0})$, there is no caustic, as expected. The caustic only occurs when $r=0$, as the coordinate curves meet, by definition.
@@
$r^{\prime}$ $=l-l_{0}-\frac{1}{\alpha v_{0}}\left[\mbox{cotgh}(\Theta-\Theta_{0})-\mbox{cotgh}\Theta_{0}\right].$ (17)
In order to connect these expression with Eqs. (10) and Eq. (11), the initial conditions must satisfy $\alpha(\tau_{0}+v_{0}l_{0})=2\Theta_{0}$ and $l_{0}=2\mbox{cotgh}\Theta_{0}/(\alpha v_{0})$. Furthermore, the contracting nature of the dS manifold can be seen by evaluating $H_{BGV}$ given by eq. (3). Considering the “observer” appearing in the definition to be the co-moving one, $u_{0}^{\mu}=(1,0,0,0)$, then
@@ IV Extension for a general flat FLRW metric
We focus on past-incomplete space-times. Evidently, if the domain of the cosmological time does not go up to minus infinity, the space-time is already incomplete, and a completion is only possible in the absence of scalar curvature singularities. Therefore, our first assumption is:
@@
Thence, the condition for the metric components, as well as its determinant, to be well-behaved when $a\rightarrow 0$ reads
@@
For the spatial $g_{ll}$ element in the plane of motion we get
@@
If assumption (iii) is not satisfied, the geometry contains a parallelly propagated curvature singularity, see Ref. YQ18 , which forbids the geometry completion.
@@ V Conclusion
Notice that, when an extension is possible, the non co-moving observer experiences a kinematic bounce exclusively due to its velocity, and hence there is no need for violation of the null energy condition: it moves in the locally contracting part reaching the $a=0$ boundary, emerging in the expanding region accessible to the co-moving observer after the bounce. This one only experiences the eternal expanding phase. Note that the spatial geometry is a curved inhomogeneous and anisotropic hyper-surface, and the bounce occurs in different proper times for each observer. Therefore, no violation of the null energy condition is necessary, for instance the dS example or a toy model with $a\propto\mbox{sech}(\alpha t)$. However, it may occur, see the discussion in YQ18 .
@@
Note that a past asymptotically dS geometry, as in Definition 3, does not necessarily have all the symmetries of a dS space-time. Possible physical realizations are scalar fields with the non-singular potentials given in Ref. YQ18 , such as the Higgs-like case. Anyway, whatever scenario one might have in mind, it seems to us quite premature to discard the extension described in this paper, and others, with arguments of implausible initial conditions. There is no sufficient knowledge of Quantum Gravity and Quantum Cosmology, which is expected to be the theory of initial conditions for Cosmology hartle , to say much about this fundamental and contrived issue.

### v2

# Bouncing completion of eternal inflation

@@ I Introduction
This was perceived by Borde, Guth and Vilenkin, who contrived it into the BGV theorem N3 ; Borde_1994 , providing a route to verify incompleteness: if the average expansion of the spatial sections is positive, the space-time is geodesically incomplete and, consequently, singular Ellis1977 . An alternative definition of singularity has been proposed in Belinski_2017 .
@@
Recently, the theorem gained renewed attention as cosmologies with cyclic properties, such as periodic Hubble functions and/or conformal cycles IS19 ; IS21 ; P10 , were argued to be incomplete KS22 ; N4 . Further, counter-examples were found due to the existence of loopholes in the construction of the average expansion LED23 , afterwards amended in a re-proposal GHZ24 . Still, past eternal inflation must be modified and, together with any eternal Universe proposal, can only be complete at the cost of breaking the null energy condition EL24a ; EL24b ; carballorubio2024 . However, a fundamental assumption in all approaches is that the space-time geometry is inextensible N2 ; Wald1984rg ; Penrose65 ; GEROCH1968526 ; Hawking66-I ; Hawking66-II ; Hawking67-III ; Hawking-Penrose ; Ellis1977 ; Senovilla_2015 . Thus, the question of whether an averaged expanding space-time can be past eternal depends on whether it admits an extension or not, which enters as an hypothesis in the BGV theorem. Along this route, extendibility conditions for singular space-times were pursued in Clarke-local-extensions ; Clarke-local-extensionsII ; Clarke-singularities-hyperbolicST , and for flat FLRW in Klein_2017 ; LP17 ; YQ18 ; HISC22 ; Nomura_2021 , through co-moving coordinates, curvature tensors in parallelly propagated tetrads and/or null-like coordinates for the case of quasi-de Sitter N5 . The aim of this paper is to propose a $\mathcal{C}^{2}$-extension of such space-times through the construction of a physically motivated chart given by the non co-moving observer reference frame. The $\mathcal{C}^{2}$ requisite for the extension comes from Einstein’s equations, which are second order coupled non-linear partial differential equations for the metric. As an intuitive argument note that, together with the infinite time dilation of the co-moving observer clock, there must be an infinite contraction of their spatial rulers, inducing a zero spatial volume in this asymptotic limit. Hence, the spatial volume seen by the non co-moving observers in this asymptotic limit may be finite, allowing its extension. Indeed, this approach allows us to exhaust all possible cases in flat FLRW space-times, showing that only past asymptotic de Sitter (dS) can be extended (excluding the cyclic models presented in Refs. IS19 ; IS21 ). There, the non co-moving observer sees his co-moving volume reaching a velocity dependent finite minimum, before growing again: the extension must contain a bounce. This paper is organized as follows: In Section II we revisit the BGV theorem, restating its assumptions, some loopholes, and the revised proposal. In section III we present a local analytic extension to non co-moving observers in the spatially flat dS space-time. In section IV we derive necessary conditions for a $\mathcal{C}^{2}$ metric extension of an arbitrary spatially flat FLRW space. We conclude in section V.
@@ II The Borde-Guth-Vilenkin Theorem
the only degree of freedom being the scale factor $a\equiv a(t)$, whose dynamics is determined by an energy-momentum tensor sourcing Einstein’s equations. Nonetheless, the following theorem and the upcoming results in this work acquire some of its strength and generality in being purely geometrical results, i.e., by making no reference to the matter content generating space-time curvature or any energy condition. Only in the end of the paper we discuss the implications of our results to the energy-momentum distribution of matter fields using General Relativity.
The geodesic equations are:
@@
$H_{BGV}(s)=\frac{-u^{\mu}_{0}}{\sigma^{2}+\kappa}\frac{Du_{\mu}}{\textrm{d}s}=\frac{-1}{\sigma^{2}+\kappa}\frac{\textrm{d}\sigma}{\textrm{d}s}=\frac{\textrm{d}}{\textrm{d}s}\mathcal{F}(\sigma),$ (3)
being $\sigma=u_{0}^{\nu}u_{\nu}$, $\kappa=u_{0}^{\nu}u_{0\nu}$, and $D/\textrm{d}s\equiv u_{0}^{\nu}\nabla_{\nu}$ is the absolute derivative along the observer’s worldline (note that the vector with the lower zero index $u_{0}^{\mu}$ is the observer measuring the expansion of the congruence given by $u^{\mu}$, with no additional indexation). For a time-like $u^{\mu}_{0}$, $\sigma=(1-v_{rel}^{2})^{-1/2}$, i.e., represents the Lorentz factor for the relative spatial velocity $v_{rel}$ between both observers, and $\kappa=-1$, while for a light-like $u^{\mu}_{0}$, $\sigma=\textrm{d}t/\textrm{d}s$, the change of time w.r.t. the affine parameter, and $\kappa=0$. Moreover,
$\mathcal{F}(\sigma)=\begin{cases}\sigma^{-1}&\text{when $\kappa=0$,}\\ \frac{1}{2}\mathrm{ln}\frac{\sigma+1}{\sigma-1}&\text{when $\kappa=-1$.}\\ \end{cases}$ (4)
To grasp some physical intuition into the significance of $\mathcal{F}$, note that, given a time-like observer $u^{\mu}_{0}$ and a non-relativistic relative velocity, $v_{rel}=|\vec{u}|\ll 1$, we find that
$\mathcal{F}\approx-\mbox{ln}\,v_{rel}.$ (5)
Hence,
$H_{BGV}\approx-\frac{\textrm{d}}{\textrm{d}s}\mbox{ln}\,v_{rel}=\frac{\Delta u_{r}}{\Delta r},$ (6)
where $\Delta r=|\Delta\vec{r}|=|\vec{u}\Delta s|$ is the separation of two test particles in the congruence $u^{\mu}$ by elapsed time $\Delta s$, and $\Delta u_{r}=\Delta\vec{u}\cdot\Delta\vec{r}/\Delta r$ is the radial projection of $\Delta\vec{u}\equiv\vec{v}_{rel}=-(\textrm{d}\vec{u}/\textrm{d}s)\Delta s$, which is the relative velocity between test particles. As a consequence, $H_{BGV}$ is just the fractional change in the velocity of the congruence $u^{\mu}$, as expected (see Ref. N3 for details).
Alongside that, the generalized Hubble parameter reduces to the usual one $H=(1/3)u^{\mu};_{\mu}$ when smeared over all directions of $u^{\mu}_{0}$ N3 . Further, defining the averaged expansion rate $\mathcal{H}_{avg}$ as
@@
$\int^{s_{f}}_{s_{i}}H_{BGV}\,\textrm{d}s=\mathcal{F}(\sigma_{f})-\mathcal{F}(\sigma_{i})\leq\mathcal{F}(\sigma_{f}).$ (8)
@@
$0<\mathcal{H}_{avg}\leq\frac{\mathcal{F}(\sigma_{f})}{\Delta s},$ (9)
@@
Nonetheless, the present work intends to scrutinize a different aspect of the BGV theorem, namely, the assumption that the space-time is inextensible. Important works have been done in this direction Klein_2017 ; LP17 ; YQ18 ; HISC22 ; Nomura_2021 . However, here we seek a different approach by constructing a chart adapted to the incomplete non co-moving observer, allowing us to grasp some physical intuition on possible extensions and the elusive nature of some singular space-times.
@@
The de Sitter space-time, a maximally symmetric solution to Einstein’s equations with positive cosmological constant N2 , is a pivotal example. The coordinate chart in which the spatial section are flat homogeneous and isotropic hypersurfaces, given by eq. (1) with an exponential scale factor $a(t)=a_{dS}(t)=\textrm{e}^{\alpha t}$, satisfies all the BGV theorem hypothesis and is therefore incomplete, albeit the absence of any singularity, indicating that the incompleteness is just a limitation of the flat patch.
@@
$r$ $=l-\frac{1}{\alpha v_{0}}\left[\mbox{coth}(\Theta-\Theta_{0})+\mbox{coth}\Theta_{0}\right],$ (13)
@@
Equations (12,13) give the coordinate transformations from the co-moving coordinates to the non co-moving ones. Direct derivation w.r.t. the parameter $l$ leads to the normalized vector field $v^{\mu}=(v^{t},v^{r},0,0)$ tangent to the space-like hyper-surfaces in the new coordinate system:
@@
where $a\equiv a_{dS}(\tau,l)$ and $r\equiv r(\tau,l)$ as in Eqs. (12) and (13). Notice that, given a value of $v_{0}$ and requiring the motion of the geodesics to occur in the $t-r$ plane, the congruence given by $u^{\mu}$ and, consequently, the proper time $\tau$, is defined and unique up to a time translation. Nonetheless, the coordinates in the surfaces orthogonal to $u^{\mu}$ are arbitrary and can be changed by different choices of parametrization.
Further, although $a_{dS}$ is expressed in terms of a hyperbolic sine, it depends on $l$ also, and the metric above is spatially inhomogeneous and anisotropic. Hence, it cannot be understood as the incomplete dS with negative spatial curvature with a zero volume element when $a=0$, as it was done in Ref. N4 , which is clearly not the case here. Note that $l$ is not a radial coordinate.
@@
which reaches a minimum $\sqrt{-g}=|v_{0}|\mbox{sin}\theta/\alpha^{2}$ when $\Theta=\Theta_{0}$, growing again for $\Theta<\Theta_{0}$. This is expected as the dS manifold can be covered by two flat patches, and if we assume that time for the non co-moving observer goes only in one direction, one sheet is contracting while the other expands. This is in contrast with Ref. N6 , in which the complete dS space-time is covered by two expanding sheets, and time runs in different directions and no flow of particles between the sheets must be imposed as a boundary condition. The argument is that the energy density of test particles traveling in the non co-moving geodesics, i.e., the $T^{0}_{0}$ component of its energy-momentum tensor, diverges in the boundary. However, this is only a coordinate problem. For the co-moving observer the non co-moving one approaches light velocity near this boundary, appearing to have infinity energy, but nothing happens from the point of view of non co-moving observers themselves, analogous to what happens near a black hole’s horizon, where observers can naturally pass through it from their own point of view.
@@
$u^{\mu}_{;\mu}=\alpha\mbox{tanh}(\Theta-\Theta_{0})+\\ 2\alpha\left[\frac{(\alpha v_{0}l+\mbox{coth}\Theta_{0})\mbox{coth}(\Theta-\Theta_{0})-1}{\alpha v_{0}l-\mbox{coth}(\Theta-\Theta_{0})+\mbox{coth}\Theta_{0}}\right].$ (17)
When $\Theta=\Theta_{0}$ the expansion becomes $\theta=2\alpha(\alpha v_{0}l+\mbox{coth}\Theta_{0})$, there is no caustic, as expected. The caustic only occurs when $r=0$, as the coordinate curves meet, by definition.
@@
$r^{\prime}$ $=l-l_{0}-\frac{1}{\alpha v_{0}}\left[\mbox{coth}(\Theta-\Theta_{0})-\mbox{coth}\Theta_{0}\right].$ (19)
In order to connect these expression with Eqs. (12) and Eq. (13), the initial conditions must satisfy $\alpha(\tau_{0}+v_{0}l_{0})=2\Theta_{0}$ and $l_{0}=2\mbox{coth}\Theta_{0}/(\alpha v_{0})$. Furthermore, the contracting nature of the dS manifold can be seen by evaluating $H_{BGV}$ given by eq. (3). Considering the “observer” appearing in the definition to be the co-moving one, $u_{0}^{\mu}=(1,0,0,0)$, then
@@ IV Extension for a general flat FLRW metric
We focus on past-incomplete space-times. Evidently, if the domain of the cosmological time does not go up to minus infinity, the space-time is already incomplete, and a completion is only possible in the absence of scalar curvature singularities. In a FLRW space-time, there are only three independent scalar quantities, namely:
$\begin{gathered}R=R^{\mu}_{\,\,\mu}=6(\dot{H}+2H^{2}),\\ K=R^{\mu\nu\lambda\rho}R_{\mu\nu\lambda\rho}=12[\dot{H}^{2}+2H^{2}(\dot{H}+H^{2})],\\ R^{\mu\nu}R_{\mu\nu}=3[(3H^{2}+2\dot{H})^{2}+3H^{4}],\end{gathered}$ (21)
the Ricci and Kretschmann scalars, and the Ricci contracted with itself, respectively. With that, our first assumption is:
@@
On the contrary, if $t\in[t_{in},t_{0}]$, with $t_{in}$ finite, either $a(t_{in})=0$ and there is a singularity, as in any Big Bang model, or $a(t_{in})=a_{in}\neq 0$ and it is trivially extensible, while if $\lim_{t\rightarrow-\infty}a(t)\neq 0$, the space-time is past complete N5 .
@@
Thence, the condition for the metric components, as well as its determinant, to be well-behaved when $a\rightarrow 0$, i.e. to be continuous and non-degenerate at these points, reads
@@
For the spatial $g_{ll}$ element we get
@@
If assumption (iii) is not satisfied, the geometry contains a parallelly propagated curvature singularity, see Ref. YQ18 , which forbids completion.
@@ V Conclusion
Notice that, when an extension is possible, the non co-moving observer experiences a kinematic bounce exclusively due to its velocity, and hence there is no need for violation of the null energy condition: it moves in the locally contracting part reaching the $a=0$ boundary, emerging in the expanding region accessible to the co-moving observer after the bounce. This one only experiences the eternal expanding phase. Note that the spatial geometry is a curved inhomogeneous and anisotropic hyper-surface, and the bounce occurs in different proper times for each observer. Therefore, no violation of the null energy condition is necessary. Some examples are de Sitter space-time itself (obviously), and the toy model with $a\propto\mbox{sech}(\alpha t)$. Of course, there are also models which are extensible according to Theorem 1, but violate the null energy condition. For example, assuming the leading term in the energy density generating the space-time curvature to be a cosmological constant, sub-leading terms with constant equation of state must violate the null energy condition in order to have a finite $\dot{H}/a^{2}$ in the past-boundary, see the discussion in YQ18 .
@@
Note that a past asymptotically dS geometry, as in Definition 3, does not necessarily have all the symmetries of a dS space-time. Possible physical realizations are scalar fields with the non-singular potentials given in Ref. YQ18 . There, conditions are given for slow-roll inflationary models, that fulfill assumptions (i) and (ii), to satisfy assumption (iii). For example, it is shown that a Starobinsky model has a p.p. curvature singularity, while a Higgs-like potential can be extended. Anyway, whatever scenario one might have in mind, it seems to us quite premature to discard the extension described in this paper, and others, with arguments of implausible initial conditions. There is no sufficient knowledge of Quantum Gravity and Quantum Cosmology, which is expected to be the theory of initial conditions for Cosmology hartle , to say much about this fundamental and contrived issue.
As generalizations of the present work, one first step should be to investigate its extension to Friedman models with curved space-like hypersurfaces. For the more involved anisotropic and non-homogeneous cases, the generalized Hubble factor $\mathcal{H}_{BGV}$, and the space-like volume of space-time using an appropriate choice of time may be useful tools to work with, but this will be the subject of future work.